# A Cognitive Process-Inspired Architecture for Subject-Agnostic Brain Visual Decoding

### v1

@@ 4.4.2 Qualitative Results
Figure 5: Compared with GLFA Li et al. (2024), the qualitative comparison demonstrates that VCFlow achieves superior semantic fidelity and temporal coherence, effectively capturing fine-grained semantics and preserving motion information in a subject-agnostic setting.
@@ 4.5 Ablation Studies
In this section, we evaluate the effectiveness of four key components: fMRI-to-image pretraining, HCAM, SARA, and HED. All experiments follow a subject-agnostic setting, where models are trained on subjects 2 and 3 and directly tested on subject 1. As shown in Table 2, each module contributes to performance gains across both video-based and image-based metrics.
Adding HCAM enhances semantic understanding, while SARA boosts cross-subject transferability, reflected in CLIP-pcc and PSNR. Introducing HED provides the most significant gains, particularly in high-level semantics and reconstruction quality.
@@ 4.6 Interpretation Results
By adopting the cortical projection technique introduced in Yang et al. (2024), we visualize the hierarchical embeddings on the brain surface map. As illustrated in Fig. 6, the early vision embeddings exhibit a clear correspondence with early visual areas, notably V1 through V4. The ventral stream embeddings show strong activation in high-level visual regions such as the FFA and PPA, while also demonstrating more diffuse projections across broader cortical areas, likely due to their encoding of abstract global semantics. In contrast, the dorsal stream embeddings align well with motion-related regions, particularly those associated with MST. These interpretation results are highly consistent with the neurocognitive structure underpinning our proposed framework.
@@ A.4 Progressive Learning Strategy
We adopt a progressive learning strategy Wang et al. (2025) to jointly optimize multiple loss functions. This strategy regulates the weight of each loss along a smooth schedule: starting from 1, increasing to 10, and then decaying back to 1 within a predefined period. The scheduling is applied to four loss functions, each with a distinct offset period to promote balanced training.
@@
$w=1+9\cdot\left|\sin\!\left(\frac{C}{T}\cdot\pi\right)\right|.$
@@ Appendix B Details of Data Pre-processing
We employ both the fMRI-to-image datasets and the fMRI-to-video dataset in our experiments Wen et al. (2018); Shen et al. (2019); Horikawa & Kamitani (2017). For fMRI data preprocessing, we first aligned all fMRI volumes to the 32k_fs_LR brain surface space based on anatomical structure Glasser et al. (2013). Unlike conventional fMRI decoding methods that flatten each fMRI frame into a one-dimensional signal and select subject-specific activated voxels, our approach transforms the fMRI data into a unified surface-based representation across subjects. After surface alignment, we applied voxel-wise $z$-transformation to normalize the fMRI signals and unfolded the cortical surface into a two-dimensional plane, thereby preserving spatial relationships between adjacent voxels. Given that only a subset of brain regions is typically activated during visual stimulation tasks Huang et al. (2021), we further restricted the analysis to early and higher visual cortical Regions of Interest (ROIs). These ROIs, encompassing a total of 8,405 vertices, were defined according to the HCP-MMP atlas Glasser et al. (2016) in the 32k_fs_LR surface space. Each processed fMRI volume was then converted into a single-channel $256\times 256$ image. For datasets with multiple runs corresponding to the same video stimulus, we averaged the aligned fMRI frames across runs. Finally, for fMRO-to-Video dataset, to account for the hemodynamic response delay—the time lag between stimulus presentation and the peak of the BOLD response—a temporal shift of approximately 6 seconds was applied to the fMRI data.
@@ C.1 Comparison with Subject-Specific Methods
We also compare our method with state-of-the-art approaches under different settings as shown in Table 4. Notably, when compared with the subject-specific state-of-the-art method NEURONS Wang et al. (2025), our approach exhibits only a modest decrease of 7% in average performance. However, it achieves the advantage of direct testing without the need for retraining.
@@
NEURONS Wang et al. (2025) Subject-specific 20.6% 81.0% 0.373 9.591 25.4% 86.2% 0.932
GLFA Li et al. (2024) Subject-adaptive pretraining 11.6% 77.7% 0.172 - 19.3% 84.7% -
@@
NEURONS Wang et al. (2025) Subject-specific 21.4% 81.0% 0.353 9.502 25.2% 86.0% 0.933
GLFA Li et al. (2024) Subject-adaptive pretraining 10.5% 76.8% 0.167 - 17.5% 83.8% -
@@
NEURONS Wang et al. (2025) Subject-specific 21.0% 81.7% 0.369 9.488 27.8% 86.8% 0.937
GLFA Li et al. (2024) Subject-adaptive pretraining 12.6% 78.1% 0.181 - 17.9% 83.7% -
@@
NEURONS Wang et al. (2025) Subject-specific 21.0% 81.2% 0.365 9.527 26.1% 86.3% 0.934
GLFA Li et al. (2024) Subject-adaptive pretraining 11.6% 77.5% 0.173 - 18.2% 84.1% -
@@
Table 4: Comparison of VCFlow with GLFA Li et al. (2024) and NEURONS Wang et al. (2025) across subjects. GLFA adopts subject-adaptive pretraining by using fMRI data from all subjects, while NEURONS is trained and evaluated on the same subject.
@@ C.2 Visualization Results
We further conduct a qualitative comparison among VCFlow, GLFA Li et al. (2024), and NEURONS Wang et al. (2025), as illustrated in Fig. 7. Compared to GLFA, VCFlow consistently delivers superior performance in both semantic fidelity and dynamic coherence. In comparison with the subject-specific model NEURONS, which exhibits high visual fidelity and strong temporal consistency, VCFlow achieves similarly high-quality reconstructions while maintaining robust motion dynamics. Notably, in scenarios involving common semantic structures, VCFlow even surpasses NEURONS by producing smoother motion trajectories and more semantically accurate content, despite being trained in a subject-agnostic manner.

### v2

@@ 4.4.2 Qualitative Results
Figure 5: Compared with GLFA (Li et al., 2024), the qualitative comparison demonstrates that VCFlow achieves superior semantic fidelity and temporal coherence, effectively capturing fine-grained semantics and preserving motion information in a subject-agnostic setting.
@@ 4.5 Ablation Studies
In this section, we evaluate the effectiveness of four key components: fMRI-to-image pretraining, HCAM, SARA, and HED. All experiments follow a subject-agnostic setting, where models are trained on subjects 2 and 3 and directly tested on subject 1. As shown in Table 2, each module contributes to performance gains across both video-based and image-based metrics. Adding HCAM enhances semantic understanding, while SARA boosts cross-subject transferability, reflected in CLIP-pcc and PSNR. Introducing HED provides the most significant gains, particularly in high-level semantics and reconstruction quality.
@@ 4.6 Interpretation Results
By adopting the cortical projection technique introduced in (Yang et al., 2024), we visualize the hierarchical embeddings on the brain surface map. As illustrated in Fig. 6, the early vision embeddings exhibit a clear correspondence with early visual areas, notably V1 through V4. The ventral stream embeddings show strong activation in high-level visual regions such as the FFA and PPA, while also demonstrating more diffuse projections across broader cortical areas, likely due to their encoding of abstract global semantics. In contrast, the dorsal stream embeddings align well with motion-related regions, particularly those associated with MST. These interpretation results are highly consistent with the neurocognitive structure underpinning our proposed framework.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by the Joint Research Scheme (JRS) under the National Natural Science Foundation of China (NSFC) and the Research Grants Council (RGC) of Hong Kong (Project No. N_HKUST654/24); the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference No. AoE/E-601/24-N and Project No. R6005-24); and the National Natural Science Foundation of China (Grant No. 62306254).
@@ A.4 Progressive Learning Strategy
We adopt a progressive learning strategy (Wang et al., 2025) to jointly optimize multiple loss functions. This strategy regulates the weight of each loss along a smooth schedule: starting from 1, increasing to 10, and then decaying back to 1 within a predefined period. The scheduling is applied to four loss functions, each with a distinct offset period to promote balanced training.
@@
$w=1+9\cdot\left|\sin\!\left(\frac{C}{T}\cdot\pi\right)\right|.$ (13)
@@
A.5 Subtasks Descriptions
Our task setup is inspired by NEURONS (Wang et al., 2025), and we describe each subtask in detail below. We first present the ventral-related tasks, which are mainly designed to capture high-level semantic information along the visual processing stream.
Concept Recognition. To enhance conceptual understanding, we introduce a concept recognition task by adding a multi-label classifier $\mathcal{D}_{cls}(\cdot)$ that predicts the key concepts in each frame from the fMRI-derived visual embeddings. Concretely, we apply a cross-entropy loss between the classifier prediction and the ground-truth (GT) concept list:
$\mathcal{L}_{cls}=\mathcal{L}_{ce}(\mathcal{D}_{cls}(\bar{e}^{v}),\mathcal{C}),$ (14)
where $\bar{e}^{v}$ denotes the mean of ${e}^{v}$ along the frame axis, and $\mathcal{C}$ is the GT concept list.
Scene Description. To further model scene-level semantics, we incorporate a scene description task that aims to generate a descriptive caption for each video frame. Specifically, we finetune a pre-trained text decoder $\mathcal{D}_{caption}(\cdot)$, which takes the fMRI-derived text embeddings $e^{t}$ as input and produces the caption. We adopt GPT-2 (Radford et al., 2019) as the text decoder and train it using prefix language modeling. Given a GT caption token sequence $\mathcal{S}=\{s_{0},s_{1},\dots,s_{|\mathcal{S}|}\}$ and the corresponding text embedding $e^{t}$, the decoder $\mathcal{D}_{caption}(\cdot)$ is trained to reconstruct $\mathcal{S}$ conditioned on $e^{t}$ as the prefix. The training objective $\mathcal{L}_{caption}$ is defined as:
$\mathcal{L}_{caption}=-\frac{1}{|\mathcal{S}|}\sum^{|\mathcal{S}|}_{i=1}\log\mathcal{D}_{caption}(s_{i}\mid s_{<i},e^{t}),$ (15)
where $s_{i}$ denotes the $i$-th token in the GT sequence $\mathcal{S}$.
Next, we consider early-vision tasks, which primarily aim to learn coarse object contours and spatial masks.
Key Object Segmentation. To better capture object-level information, we design a text-driven video decoder based on the VAE video decoder (von Platen et al., 2022). This decoder takes the video embeddings $e^{v}$ together with the text embeddings $e^{t}$ as inputs. For this task, the text embeddings are obtained by encoding the category name of the key object with the CLIP text encoder. A cross-attention module is then applied to activate specific patches in $e^{v}$ (used as queries $Q$) corresponding to $e^{t}$ (used as keys $K$ and values $V$): ${e}^{seg}=\mathrm{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d}}\right)\cdot V.$ The activated feature ${e}^{seg}$ is upsampled to a higher resolution for pixel-level prediction. We then employ a simple segmentation head $\mathcal{D}_{vs}(\cdot)$ to generate the binary segmentation masks $y_{seg}$ for the key object in the video. The training objective for this task is a binary cross-entropy loss, denoted as $\mathcal{L}_{seg}$.
Finally, we introduce the dorsal-related task, which focuses on modeling global motion information.
Blurry Video Reconstruction. We reuse the same VAE decoder as in the key object segmentation task, but replace the segmentation head $\mathcal{D}_{vs}(\cdot)$ with a reconstruction head $\mathcal{D}_{vr}(\cdot)$. Given the video embeddings $e^{v}$, the text-driven video decoder together with $\mathcal{D}_{vr}(\cdot)$ generates a blurry video $y^{motion}_{c}$. We then map $y^{motion}_{c}$ into the latent space of the Stable Diffusion VAE to obtain the latent embeddings $y^{\prime}_{c}$. This subtask is trained with a mean absolute error (MAE) loss, defined as:
$\mathcal{L}_{motion}=\frac{1}{F}\sum^{F}_{i=1}\big\lvert y^{motion}_{c,i}-y^{\prime}_{c,i}\big\rvert$. (16)
It is worth noting that, unlike NEURONS, which jointly optimizes all tasks over a single global feature representation, we explicitly associate each task with a particular cognitive process and train it on the corresponding process-specific features.
A.6 Inference
At inference time, our pipeline leverages a pre-trained T2V diffusion model (Guo et al., 2023), in a setup similar to NEURONS (Wang et al., 2025). We conditions the model jointly on a control image, a blurry video, and a text description. These three inputs are assembled from the outputs of the decoupled tasks: a control image is reconstructed from each frame via unCLIP (Ramesh et al., 2022), while $\mathcal{D}_{cls}$ and $\mathcal{D}_{caption}$ provide the predicted concepts and generated captions, respectively. The embedding of the top-1 predicted concept, together with the caption embedding, is used to guide video-mask prediction through $\mathcal{D}_{vs}$ and blurry video reconstruction through $\mathcal{D}_{vr}$. To further enhance the prominence of the key object, we rescale its binary mask from 0,1{0,1} to $[0.5,1]$ and multiply it with both the control image and the blurry video before feeding them into the diffusion model.
Because inference for unseen subjects does not require any subject-specific finetuning, their data can be directly processed by the model. With this setup, the inference time—normalized by the total number of videos—remains below 10 seconds per video. All measurements are obtained on an NVIDIA RTX 4090 GPU.
A.7 Details of Brain Model
As shown in Fig. 3, our Brain Model extracts four types of representations from the input fMRI signals: a full-brain representation $\mathbf{E}_{\text{brain}}$, early visual features $\mathbf{E}_{\text{early}}$, ventral-stream features $\mathbf{E}_{\text{ventral}}$, and dorsal-stream features $\mathbf{E}_{\text{dorsal}}$. Here, the Brain Model refers to the module responsible for deriving these fMRI-based features, consisting of a ViT backbone together with several linear projection heads.
[…]
@@ Appendix B Details of Data Pre-processing
We employ both the fMRI-to-image datasets and the fMRI-to-video dataset in our experiments (Wen et al., 2018; Shen et al., 2019; Horikawa and Kamitani, 2017). For fMRI data preprocessing, we first aligned all fMRI volumes to the 32k_fs_LR brain surface space based on anatomical structure (Glasser et al., 2013). Unlike conventional fMRI decoding methods that flatten each fMRI frame into a one-dimensional signal and select subject-specific activated voxels, our approach transforms the fMRI data into a unified surface-based representation across subjects. After surface alignment, we applied voxel-wise $z$-transformation to normalize the fMRI signals and unfolded the cortical surface into a two-dimensional plane, thereby preserving spatial relationships between adjacent voxels. Given that only a subset of brain regions is typically activated during visual stimulation tasks (Huang et al., 2021), we further restricted the analysis to early and higher visual cortical Regions of Interest (ROIs). These ROIs, encompassing a total of 8,405 vertices, were defined according to the HCP-MMP atlas (Glasser et al., 2016) in the 32k_fs_LR surface space. Each processed fMRI volume was then converted into a single-channel $256\times 256$ image. For datasets with multiple runs corresponding to the same video stimulus, we averaged the aligned fMRI frames across runs. Finally, for fMRO-to-Video dataset, to account for the hemodynamic response delay—the time lag between stimulus presentation and the peak of the BOLD response—a temporal shift of approximately 6 seconds was applied to the fMRI data.
@@ C.1 Comparison with Subject-Specific Methods
We also compare our method with state-of-the-art approaches under different settings as shown in Table 4. Notably, when compared with the subject-specific state-of-the-art method NEURONS (Wang et al., 2025), our approach exhibits only a modest decrease of 7% in average performance. However, it achieves the advantage of direct testing without the need for retraining.
@@
NEURONS (Wang et al., 2025) Subject-specific 20.6% 81.0% 0.373 9.591 25.4% 86.2% 0.932
GLFA (Li et al., 2024) Subject-adaptive pretraining 11.6% 77.7% 0.172 - 19.3% 84.7% -
@@
NEURONS (Wang et al., 2025) Subject-specific 21.4% 81.0% 0.353 9.502 25.2% 86.0% 0.933
GLFA (Li et al., 2024) Subject-adaptive pretraining 10.5% 76.8% 0.167 - 17.5% 83.8% -
@@
NEURONS (Wang et al., 2025) Subject-specific 21.0% 81.7% 0.369 9.488 27.8% 86.8% 0.937
GLFA (Li et al., 2024) Subject-adaptive pretraining 12.6% 78.1% 0.181 - 17.9% 83.7% -
@@
NEURONS (Wang et al., 2025) Subject-specific 21.0% 81.2% 0.365 9.527 26.1% 86.3% 0.934
GLFA (Li et al., 2024) Subject-adaptive pretraining 11.6% 77.5% 0.173 - 18.2% 84.1% -
@@
Table 4: Comparison of VCFlow with GLFA (Li et al., 2024) and NEURONS (Wang et al., 2025) across subjects. GLFA adopts subject-adaptive pretraining by using fMRI data from all subjects, while NEURONS is trained and evaluated on the same subject.
@@ C.2 Visualization Results
We further conduct a qualitative comparison among VCFlow, GLFA (Li et al., 2024), and NEURONS (Wang et al., 2025), as illustrated in Fig. 7. Compared to GLFA, VCFlow consistently delivers superior performance in both semantic fidelity and dynamic coherence. In comparison with the subject-specific model NEURONS, which exhibits high visual fidelity and strong temporal consistency, VCFlow achieves similarly high-quality reconstructions while maintaining robust motion dynamics. Notably, in scenarios involving common semantic structures, VCFlow even surpasses NEURONS by producing smoother motion trajectories and more semantically accurate content, despite being trained in a subject-agnostic manner.
@@ C.3 Failure Cases
C.4 Additional Ablation Results
To strengthen the persuasiveness of our SARA and HED designs, we conducted ablation studies on both components, as shown in Table 5 and Table 6.
HCAM SARA HED Frame-based Video-based
[…]
7.52% 68.2% 0.392 9.090 9.67% 77.3% 0.903
10.7% 75.0% 0.368 9.983 13.9% 81.7% 0.924
14.2% 78.6% 0.389 10.469 18.9% 84.8% 0.944
Table 5: Ablations on the components of SARA, and all results are from subject 1.
HCAM SARA HED Frame-based Video-based
[…]
10.0% 72.6% 0.356 9.370 12.8% 81.4% 0.907
8.2% 71.1% 0.360 10.197 14.6% 82.3% 0.970
13.2% 78.9% 0.408 10.737 16.4% 84.1% 0.942
11.3% 76.0% 0.383 10.123 15.2% 82.4% 0.926
14.2% 78.6% 0.389 10.469 18.9% 84.8% 0.944
Table 6: Ablations on the components of HED, and all results are from subject 1.
C.5 Visualization Comparison
We conducted a comparative visualization of the preprocessed data and the semantic tokens, as shown in Fig. 9.
Figure 9: Visualization of the preprocessed data and the semantic tokens.